OpenStudy (anonymous):

if x-y=101 and x+y=135 what is x and y?

OpenStudy (anonymous):

x-y = 101 x+y= 135 2x = 236 x=118 118-y = 101 Note: y = 17 119+y = 135 Note: y = 17 ans: x = 118 and y = 17

OpenStudy (anonymous):

awesome thanks